OpenStudy (anonymous):

-3 and 1/3 • 5 and 1/4

OpenStudy (aripotta):

convert them into improper fractions and then multiply

OpenStudy (lizz123):

\[3\frac{ 1 }{ 3 }\] Multiply \[3\times3+1=?\] What does that equal??

OpenStudy (lizz123):

@queenofnothing

OpenStudy (anonymous):

10 sorry for some reason it wont let me reply on my phone.

OpenStudy (lizz123):

It is -10.

OpenStudy (lizz123):

So now \[5 \frac{ 1 }{ 4 }\] Do the same. Multiply and then add \[4\times5+1\]

OpenStudy (lizz123):

I meant -3 * 3 + 1 = -8

OpenStudy (anonymous):

ok 21.

OpenStudy (lizz123):

Yeah so rewrite that

OpenStudy (lizz123):

yeah it equals 21. So it now goes into 21/4

OpenStudy (lizz123):

For -8. It turns -8/3

OpenStudy (lizz123):

\[\frac{ 21 }{ 4 }\times \frac{ -8 }{3 }\] Multiply 21 and -8. What does that equal

OpenStudy (anonymous):

-168

OpenStudy (lizz123):

Yeah so now you have\[\frac{ -168 }{ 4\times3 }\] Multiply 4*3

OpenStudy (lizz123):

@queenofnothing

OpenStudy (anonymous):

12

OpenStudy (lizz123):

yeah \[\frac{ -168 }{ 12 }\] Now put it into a mixed number

OpenStudy (lizz123):

Which the answer would be -12 0/14